Solve by the elimination method.
x+8y=29
-x+3y=4
I do not understand this 'elimination method', can someone please help me??

Answer by oscargut(2103)

Solve by the elimination method.
x+8y=29
-x+3y=4
you have to multiply any equation for a number to eliminate x or y after adding the equations.
in this example you do not have to do nothing before the addition because
if you do it:
11y=33 (you eliminate the x)
then y = 3
after that you substitute
and
x+8(3)=29
then
x=5
Answer: solution is x=5,y=3